OpenStudy (fanduekisses):

Law of Gravitation: Earth and the Moon are attracted to each other due to gravity. What would happen to the strength of the gravitational attraction between Earth and the Moon if the distance between them were doubled?

OpenStudy (fanduekisses):

Wouldn't the gravitational force be halfed?

OpenStudy (fanduekisses):

because if the distance between two objects increases, the GF decreases and vice versa.

OpenStudy (anonymous):

hm

OpenStudy (anonymous):

u know the relation ??

OpenStudy (anonymous):

like in FORMULA??

OpenStudy (anonymous):

ya....rt

OpenStudy (anonymous):

U use the formula and find it!!simple

OpenStudy (anonymous):

LAW OF GRAVITATION

OpenStudy (anonymous):

no need for numbers u shud check for ratios

OpenStudy (fanduekisses):

how?

OpenStudy (anonymous):

U have the formula with u now??

OpenStudy (fanduekisses):

Fg = (m1 x m2)/d^2

OpenStudy (anonymous):

yes ignore m1 and m2

OpenStudy (anonymous):

since masses are constant

OpenStudy (anonymous):

then we get

OpenStudy (fanduekisses):

yep

OpenStudy (anonymous):

f is proportional to \[d^2\]

OpenStudy (anonymous):

*inversely proportional

OpenStudy (anonymous):

\(\sf \Large Fg= \frac{Gm_1m_2}{d^2}\\Fg'=\frac{Gm_1m_2}{(2d)^2}\\Fg'=1/4 Fg\)

OpenStudy (anonymous):

it will not be halved, it will be 1/4 of the original Fg

OpenStudy (anonymous):

\[f(1)/f(2)=d(2)^2/d(1)^2\]

OpenStudy (anonymous):

where f(1) represents original and f(2) represents modified values

OpenStudy (anonymous):

shouldn't it be d(1)^2/d(2)^2 if f(1)/f(2) ?

OpenStudy (fanduekisses):

Ok, I understand this now, idk what happens I tend to forget things over holiday breaks lol. thanks guys.

OpenStudy (anonymous):

no problem :)